OpenStudy (kj4uts):

If your starting salary were $35,000 and you received a 6% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first 16 years that you worked? Round your answer to the nearest whole dollar, and express your answer without using commas. The first 15 years you made a total of 35k * ( 1.06^15 - 1)/(1.06 -1) = $814658.9459. If you work 16 years and assuming that you had again the 6% increase in the 16th year, then you have a total of 35k * ( 1.06^16 - 1)/(1.06 -1) =$898538

OpenStudy (kj4uts):

Im confused based on the question is the answer $814659 or $898538?

OpenStudy (confluxepic):

I already told you about 2 hours ago.

OpenStudy (confluxepic):

I got $814659.

OpenStudy (mathstudent55):

You are asked what the sum of 16 years of salary is.

OpenStudy (kj4uts):

Im trying to get other opinion because people keep telling me $814659 others are saying $898538 so who do I believe

OpenStudy (tkhunny):

If r = 1.06 35000(1 + r + r^2 + ... r^15) = 35000\(\dfrac{1-r^{16}}{1-r}\) There are 16 terms in those brackets. Getting another raise at the end of year 16 is not significant. You don't work that year in this scenario.

OpenStudy (mathstudent55):

If you work ___16___ years and assuming that you had again the 6% increase in the 16th year, then you have a total of 35k * ( 1.06^16 - 1)/(1.06 -1) =$898538

OpenStudy (kj4uts):

well what do you think based on the question that its asking

OpenStudy (mathstudent55):

The first number corresponds to 15 years of work. The second number corresponds to 16 years of work. You are asked about 16 years, not 15.

OpenStudy (tkhunny):

Asked and answered. Try this question. If your starting salary were $35,000 and you received a 6% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first year that you worked?

OpenStudy (kj4uts):

ok i don't know what this guy just put but ok. I see if it was just 15 years it would be 814659 but since its asking how much you would have made over 16 years that would be 898538

OpenStudy (confluxepic):

I miscalculated. Sorry.

OpenStudy (kj4uts):

Not a problem thank you for helping I didn't mean to put this on again I was just really confused. Thanks again though.